# On the Holographic Geometry of Deterministic Computation

### v1

@@ Abstract
Standard simulations of Turing machines suggest a linear relationship between the temporal duration $t$ of a run and the amount of information that must be stored by known simulations to certify, verify, or regenerate the configuration at time $t$. For deterministic multitape Turing machines over a fixed finite alphabet, this apparent linear dependence is not intrinsic: any length-$t$ run can be simulated in space $O(\sqrt{t})$ via a Height Compression Theorem for succinct computation trees together with an Algebraic Replay Engine. In this paper we recast that construction in geometric and information-theoretic language. We interpret the execution trace as a spacetime dependency DAG and exhibit a family of recursively defined holographic boundary summaries such that, along the square-root-space simulation, the total description length of all boundary data stored at any time is $O(\sqrt{t})$. Using Kolmogorov complexity, we prove that every internal configuration has constant conditional description complexity given the appropriate boundary summary and time index, establishing that the spacetime bulk carries no additional algorithmic information beyond its boundary. We express this as a one-dimensional computational area law: there exists a simulation in which the information capacity of the active “holographic screen” needed to generate a spacetime region of volume $t$ is bounded by $O(\sqrt{t})$. In this precise sense, deterministic computation on a one-dimensional work tape admits a holographic representation, with the bulk history algebraically determined by data residing on a lower-dimensional boundary screen.
@@ 2.2 Block-respecting runs and time-blocks
The Height Compression paper [6] shows that for suitably chosen $b$ (as a function of $t$) and after an appropriate preprocessing step, one may assume without loss of generality that a long run is block-respecting in the above sense. We treat that reduction as a black box in this paper for brevity; the full proof details are provided in the companion work.
@@ 3.1 The spacetime bulk and block decomposition
Let $\mathcal{M}$ denote the spacetime DAG of a length-$t$ run of $M$; its vertices are local degrees of freedom (tape cells and head states) at times $\tau\in\{0,\dots,t\}$, and its directed edges represent causal dependencies induced by $\delta$ between successive configurations.
Fix a block size $b$ and let $T=\lceil t/b\rceil$. The block decomposition of $\mathcal{M}$ is the collection
@@
the number of vertices in $\mathcal{M}$ (equivalently, the number of local update operations). In the one-dimensional work-tape model considered in Theorem 4, we have $V_{t}=\Theta(t)$, so bounds stated as functions of $t$ can equivalently be interpreted as bounds in terms of $V_{t}$ up to constant factors.
@@
The height-compressed computation tree $\mathcal{T}$ of Theorem 1 induces a natural hierarchical structure on these blocks. The following corollary makes explicit the “radial” behavior of block sizes along root-to-leaf paths.
@@ 3.3 Active holographic screen
and when $b$ is chosen on the order of $\sqrt{t}$, the $\log T$ term is negligible compared to $\sqrt{t}$. In particular, we obtain the one-dimensional computational area law of Theorem 4 in Section 4.
@@ 4 Bulk Vacuum and Computational Area Law
We now make precise two consequences of the structure described in Sections 2 and 3. First, we formalize the sense in which the bulk of a deterministic computation carries no independent information beyond its boundary summaries. Second, we restate the square-root space bound as a one-dimensional computational area law, and we formulate a conjectural extension to higher-dimensional local architectures.
@@ 4.1 Bulk configurations have $O(1)$ conditional complexity
Lemma 1 formalizes the following informal intuition. Once the machine $M$ and the holographic boundary state $\partial\Omega=\sigma([L,R])$ are fixed, every internal configuration $C_{\tau}$ with $\tau\in(L,R)$ is generated by a fixed, constant-complexity algebraic procedure (namely the ARE). In this sense, the bulk of the spacetime DAG $\mathcal{M}$ carries no additional algorithmic information beyond what is already present in the boundary summaries and the global description of $M$: all nontrivial information required to specify the run is concentrated on its boundary summaries, and the interior is a deterministic evaluation trace of a boundary-defined circuit.
@@ A.2 Encodings of configurations, summaries, and indices
    by concatenating encodings of $q_{\mathrm{in}}$, $q_{\mathrm{out}}$, the head positions, and the interface window $W_{\mathrm{interface}}$, again via a fixed self-delimiting scheme.
@@ C.1 Proof of Lemma 1
C.2 Proof of Theorem 4

### v2

@@ Abstract
Standard simulations of Turing machines suggest a linear relationship between the temporal duration $t$ of a run and the amount of information that must be stored by known simulations to certify, verify, or regenerate the configuration at time $t$. For deterministic multitape Turing machines over a fixed finite alphabet, this apparent linear dependence is not intrinsic: any length-$t$ run can be simulated using $O(\sqrt{t})$ work-tape cells via a Height Compression Theorem for succinct computation trees together with an Algebraic Replay Engine. In this paper we recast that construction in geometric and information-theoretic language. We interpret the execution trace as a spacetime DAG of local update events and exhibit a family of recursively defined holographic boundary summaries such that, along the square-root-space simulation, the total description length of all boundary data stored at any time is $O(\sqrt{t})$. Using Kolmogorov complexity, we prove that every internal configuration has constant conditional description complexity given the appropriate boundary summary and time index, establishing that the spacetime bulk carries no additional algorithmic information beyond its boundary. We express this as a one-dimensional computational area law: there exists a simulation in which the information capacity of the active “holographic screen” needed to generate a spacetime region of volume proportional to $t$ is bounded by $O(\sqrt{t})$. In this precise sense, deterministic computation on a one-dimensional work tape admits a holographic representation, with the bulk history algebraically determined by data residing on a lower-dimensional boundary screen.
@@ 2.2 Block-respecting runs and time-blocks
The Height Compression paper [6] shows that for suitably chosen $b$ (as a function of $t$) and after an appropriate preprocessing step, one may assume without loss of generality that a long run is block-respecting in the above sense. We treat that reduction as a black box in this paper for brevity (the full proof details are provided in the companion work). This black-box reduction ensures that the heads do not stray too far within a block and that any long excursions are folded into a structured pattern which can be summarized via a fixed-size interface window.
@@ 3.1 The spacetime bulk and block decomposition
Fix a deterministic multitape Turing machine $M$ with $k$ work tapes and Lipschitz-local head motion as in Section 2.1. Consider a length-$t$ run
$C_{0},C_{1},\dots,C_{t}$
of $M$ on some input.
The spacetime DAG of this run is a finite directed acyclic graph
$\mathcal{M}=(V,E)$
defined as follows. For each time $\tau\in\{0,\dots,t-1\}$ and each tape head $i\in\{1,\dots,k\}$, we create a vertex representing the local update neighborhood around that head and the control state at time $\tau$; thus, for every $\tau$ we introduce $O(1)$ vertices, with the implied constant depending only on $M$. Directed edges in $E$ connect vertices at time $\tau$ to vertices at time $\tau+1$ whenever the transition function $\delta$ allows information to flow between the corresponding local neighborhoods in one step.
Fix now a block size $b\in\mathbb{N}$ and let $T=\lceil t/b\rceil$. The block decomposition of $\mathcal{M}$ is the collection
@@
the number of vertices in $\mathcal{M}$ (equivalently, the number of local update events). For a fixed machine $M$ with a fixed number of tapes, there are $\Theta(1)$ such update vertices per time step, so in the one-dimensional work-tape model considered in Theorem 4 we have
$V_{t}\;=\;\Theta(t).$
Consequently, bounds stated as functions of $t$ can equivalently be interpreted as bounds in terms of $V_{t}$ up to constant factors.
@@
The height-compressed computation tree $\mathcal{T}$ of Theorem 1 induces a natural hierarchical structure on these blocks. The following remark makes explicit the “radial” behavior of block sizes along root-to-leaf paths.
@@ 3.3 Active holographic screen
and when $b$ is chosen on the order of $\sqrt{t}$, the $\log T$ term is negligible compared to $\sqrt{t}$. In particular, we obtain the one-dimensional computational area law of Theorem 4 in Section 4, where we also show that, relative to the boundary data stored on this screen, both individual bulk configurations and the entire spacetime history admit $O(1)$ conditional description complexity.
@@ 4 Bulk Vacuum and Computational Area Law
We now make precise two consequences of the structure described in Sections 2 and 3. First, we formalize the sense in which the bulk of a deterministic computation carries no independent information beyond its boundary summaries, both at the level of individual configurations and at the level of entire spacetime slabs. Second, we restate the square-root space bound as a one-dimensional computational area law, and we formulate a conjectural extension to higher-dimensional local architectures.
@@ 4.1 Bulk configurations have $O(1)$ conditional complexity
The lemma is a pointwise statement: each individual bulk configuration has constant conditional description complexity given the appropriate boundary summary and time index. The next corollary upgrades this to a whole-interval statement.
Corollary 1 (Holographic compression of interval histories).
Let $[L,R]$ be any block-respecting interval and let
$H_{[L,R]}:=(C_{L},\dots,C_{R})$
denote the corresponding configuration history, encoded as a single binary string as in Appendix A. Then there exists a constant $c^{\prime}\geq 1$, depending only on $M$, $U$, and the fixed ARE implementation, such that
$K\bigl(H_{[L,R]}\,\big|\,\sigma([L,R])\bigr)\;\leq\;c^{\prime}.$
In particular, for the full run interval (denoted here by $[0,t]$), the entire history $H_{t}:=(C_{0},\dots,C_{t})$ satisfies
$K\bigl(H_{t}\,\big|\,\sigma([0,t])\bigr)\;=\;O(1).$
Proof.
See Appendix C, Proof of Corollary 1. ∎
@@
Lemma 1 formalizes the following informal intuition. Once the machine $M$ and the holographic boundary state $\partial\Omega=\sigma([L,R])$ are fixed, every internal configuration $C_{\tau}$ with $\tau\in(L,R)$ is generated by a fixed, constant-complexity algebraic procedure (namely the ARE). In this sense, the bulk of the spacetime DAG $\mathcal{M}$ carries no additional algorithmic information beyond what is already present in the boundary summaries and the global description of $M$: all nontrivial information required to specify any single configuration is concentrated on its boundary data.
Corollary 1 upgrades this from a pointwise to a global statement: for any block-respecting interval $[L,R]$, the joint configuration sequence $H_{[L,R]}$ has $O(1)$ conditional Kolmogorov complexity given its boundary summary $\sigma([L,R])$. In particular, for the full interval $[0,t]$, the complete bulk history $H_{t}=(C_{0},\dots,C_{t})$ is, up to an additive constant in description length, an algebraically determined evaluation trace of a boundary-defined circuit.
@@ A.2 Encodings of configurations, summaries, and indices
    by concatenating encodings of $q_{\mathrm{in}}$, $q_{\mathrm{out}}$, the head positions, the interface window $W_{\mathrm{interface}}$, and, for convenience in the Kolmogorov complexity arguments, a self-delimiting encoding of the interval endpoints $(L,R)$ (or equivalently of $L$ together with the length RL+1R-L+1). This augmentation does not affect any of the combinatorial properties of interval summaries used in the main text.
@@
For an interval $[L,R]$, the history
$H_{[L,R]}:=(C_{L},\dots,C_{R})$
is encoded as a single binary string $\mathrm{enc}(H_{[L,R]})$ using any fixed self-delimiting encoding of finite sequences (for example, by prefixing each $\mathrm{enc}(C_{\tau})$ with its length in unary or via a standard pairing function). The precise choice of sequence encoding is immaterial as long as it is computable and injective.
@@ C.1 Proof of Lemma 1
C.2 Proof of Corollary 1
Recall that for any block-respecting interval $[L,R]$, we write
$H_{[L,R]}:=(C_{L},\dots,C_{R})$
for the corresponding configuration history, and we view $H_{[L,R]}$ as a single binary string $\mathrm{enc}(H_{[L,R]})$ obtained by concatenating the encodings $\mathrm{enc}(C_{\tau})$ for $\tau=L,\dots,R$ in a fixed, self-delimiting way as described in Appendix A. The statement of Corollary 1 is that there exists a constant $c^{\prime}$ such that
$K\bigl(H_{[L,R]}\,\big|\,\sigma([L,R])\bigr)\;\leq\;c^{\prime}$
for all block-respecting intervals $[L,R]$, with $c^{\prime}$ depending only on $M$, $U$, and the fixed ARE implementation.
Proof of Corollary 1.
Fix the universal prefix machine $U$, the deterministic multitape machine $M$, and the specific implementation of the Algebraic Replay Engine (ARE) from Theorem 2, exactly as in the proof of Lemma 1. Let $A$ denote the underlying Turing machine implementation of the ARE.
We define a single prefix-free program $p^{\dagger}$ for $U$ as follows.
On input $x$, interpret $x$ as $\mathrm{enc}(\sigma([L,R]))$ for some block-respecting interval $[L,R]$. Decode the interval endpoints $(L,R)$ from this encoding (recall that by convention they are included in $\sigma([L,R])$; see Appendix A). Initialize an output buffer to the empty string. For each $\tau$ in the integer range $\{L,\dots,R\}$, do:
  1. 1.
    Simulate $A$ on input $(\sigma([L,R]),\tau)$.
  2. 2.
    Let $C_{\tau}$ be the configuration output by $A$, and append $\mathrm{enc}(C_{\tau})$ to the output buffer using the fixed, self-delimiting encoding scheme for sequences described in Appendix A.
When the loop terminates, output the contents of the buffer, which is exactly the encoding of $H_{[L,R]}$.
The description (binary code) of $p^{\dagger}$ consists of:
  • a description of the fixed ARE implementation $A$, and
  • a constant amount of wrapper logic instructing $U$ to decode $\sigma([L,R])$, recover $(L,R)$, iterate over $\tau=L,\dots,R$, invoke $A$ on each $\tau$, and concatenate the resulting encodings into the final output string.
Both components are independent of $[L,R]$, of the particular run, and of the specific boundary summary; they depend only on $M$, $U$, and the chosen implementation of $A$.
Let $\ell^{\dagger}$ be the length of $p^{\dagger}$ in bits. Then, for every block-respecting interval $[L,R]$ and its interval summary $\sigma([L,R])$, when $U$ is run on program $p^{\dagger}$ with conditional input $\mathrm{enc}(\sigma([L,R]))$, it outputs $\mathrm{enc}(H_{[L,R]})$. Therefore the conditional Kolmogorov complexity of $H_{[L,R]}$ given $\sigma([L,R])$ satisfies
$K\bigl(H_{[L,R]}\,\big|\,\sigma([L,R])\bigr)\;=\;K\bigl(\mathrm{enc}(H_{[L,R]})\,\big|\,\mathrm{enc}(\sigma([L,R]))\bigr)\;\leq\;\ell^{\dagger}+O(1),$
where the $O(1)$ term is the universal overhead for using $p^{\dagger}$ as a program on $U$. Setting $c^{\prime}:=\ell^{\dagger}+O(1)$ yields the desired bound, with $c^{\prime}$ depending only on $M$, $U$, and the fixed ARE implementation.
Specializing to the full run interval $[0,t]$, we obtain $H_{[0,t]}=H_{t}$ and $\sigma([0,t])$ as in Section 4, so
$K\bigl(H_{t}\,\big|\,\sigma([0,t])\bigr)\;\leq\;c^{\prime},$
as claimed in the main text. This completes the proof of Corollary 1. ∎
C.3 Proof of Theorem 4